TEMPERATURE DEPENDENCE OF THE DIPOLE-DIPOLE COLLISION CROSS SECTION OF ^{13}CH_{3}F

\underline{LIAM DUFFY}, Department of Physics, Duke University, Durham, NC 27706; DAVID SKATRUD, Army Research Office, P.O. Box 12211, RTP, NC, 27709.

For molecules with permanent electric dipole moments, the most rapid energy transfer process is due to dipole-dipole collisions which induce ΔJ = ±1 transitions. Although this process is the dominant contributor to pressure broadening, little is known about the temperature dependence of the process. We have used a time resolved pump/probe technique based on infrared-millimeter/submillimeter wave double resonance spectroscopy to measure the temperature dependence of the dipole-dipole and higher-order multipole cross sections in ^{13}CH_{3}F. The measurements were performed over a temperature range of 200K to 400K. The experimental results are compared with the temperature scaling predicted by Anderson, Tsao and Curnette theory and the statistical power gap law.